# Adaptive Layerwise Perturbation: Unifying Off-Policy Corrections for LLM RL

### 19 Mar 2026 v1

@@ Abstract
Off-policy problems such as policy staleness and training–inference mismatch, has become a major bottleneck for training stability and further exploration for LLM RL. To enhance inference efficiency, the distribution gap between the inference and updated policy grows, leading to heavy-tailed importance ratios. Heavy-tailed ratios arise when the policy is locally sharp, which further inflates sharp gradients and can push updates outside the trust region. To address this, we propose Adaptive Layerwise Perturbation (ALP) by injecting small learnable perturbations into input hidden states of each layer during updates, which is used as the numerator of the importance ratio against the unchanged inference policy in the objective. Intuitively, by adding controlled noise to intermediate representations, ALP prevents the updated policy from deviating too sharply from the inference policy, and enlarges the policy family to cover the inference policy family with mismatch noises. Hence, the flattened distribution can naturally tighten the updated and inference policy gap and reduce the tail of importance ratios, thus maintaining training stability. This is further validated empirically. Experiments on single-turn math and multi-turn tool-integrated reasoning tasks show that ALP not only improves final performance, but also avoid blow up of importance ratio tail and KL spikes during iterative training, along with boosted exploration. Ablations show that representation-level perturbations across all layers are most effective, substantially outperforming partial-layer and logits-only variants.
Figure 1: Left: Visualization of Adaptive Layerwise Perturbation (ALP), where layerwise small perturbation variable is added to the model during training to cover the layerwise bias between training and inference policies. Right: the comparison between full training with perturbation (ALP) or without perturbation (Bypass) across $1840$ training steps. Without perturbation, the policy distribution becomes more spiky and brittle, leading to blow-up tails. In contrast, ALP smooths policies, tightens the envelope and stabilizes the importance ratio.
@@ 1 Introduction
A prominent line of work stabilizes training by modifying importance ratios. One approach changes the aggregation way of importance ratios in the objective. It replaces the ratio at each token index in GRPO (Shao et al., 2024) with the multiplication of ratios in the whole trajectory (Zheng et al., 2025), but this does not consider training-inference mismatch. Another approach targets on training-inference mismatch. For example, Bypass is a baseline mentioned in (Jiacai et al., 2025; Yao et al., 2025) that uses the ratio of the updated and rollout policies. TIS/MIS (Yao et al., 2025; Jiacai et al., 2025) further correct mismatch by multiplying an auxiliary ratio of the proxy policy (an old policy acting as the anchor) and the rollout policy. Then, they truncate or mask tokens whose auxiliary ratio exceeds a threshold. Although effective at preventing catastrophic collapse, these methods introduce practical limitations: they split off-policy effects into two separately truncated ratios, which can over-truncate updates that are otherwise within a valid trust region, thereby inducing additional bias and leading to early plateau. We defer a detailed discussion of related work to Section 2.
@@
We argue that off-policy instability in LLM RL is not solely a bookkeeping issue of “which importance ratio to use,” but also a geometry issue: the noisy update may push the policy toward sharp, brittle regions where small distribution shifts (from staleness or inference mismatch) cause large changes in action probabilities. This motivates a perturbation-sampling perspective from model smoothness and distributional learning approaches (Shen & Meinshausen, 2023; Hao et al., 2025): when the dominant failure mode is system-induced noise and staleness, a natural defense is fighting noise with structured noise.
@@
Specifically, we propose Adaptive Layerwise Perturbation (ALP) by injecting a learnable Gaussian perturbation $\delta\sim\mathcal{N}(0,\sigma^{2}I)$  [^1] The perturbation distribution is not restricted to Gaussian, it only needs to have zero mean. into the input hidden states across layers during policy updates and producing a perturbed training distribution $\pi^{\mathrm{train}}_{\theta,\sigma}$ (Figure 1: Left). We then optimize the objective using a single unified ratio of $\pi^{\mathrm{train}}_{\theta,\sigma}$ and the unperturbed inference policy i.e., perturbation is applied only to the numerator during training.
@@
Intuitively, ALP can smooth the local optimization landscape to suppress heavy-tailed ratio excursions, and enlarge the training distribution family. Moreover, this is testified by the full-training comparison in Figure 1 (Right). Starting from the same base model, the training policy without perturbation (Bypass) gradually gets sharp and brittle, and the importance ratio tail finally explode. In contrast, the training policy with perturbation (ALP) stays stable and controlling the mismatch between the training and inference system, thus tightening the ratio quantile envelop. Beyond stability, by enlarging effective support and preventing premature concentration on brittle modes, ALP also encourages exploration, particularly in multi-turn settings where compounding errors can reduce coverage.
@@
    Experiments: stability, performance, and ablations. Across single-turn and multi-turn agentic reasoning tasks, ALP improves stability and consistently outperforms MIS and Bypass. We further show that ALP reduces off-policy mismatch and exhibits more stable optimization dynamics, consistent with improved exploration and robustness. Ablations further reveal that perturbing all layers is most effective, and in partial settings, perturbations closer to lower layers tend to perform better than those restricted to upper layers.
2 Related Works
Off-policy
Recent systems efforts aim to eliminate training-inference mismatch by enforcing bitwise-consistent inference, enabling closer-to on-policy RL in practice (He & Lab, 2025) or using higher precisions (Qi et al., 2025). These directions are promising, but maintaining strict bitwise consistency is computationally inefficient, and can be challenging under common production constraints such as changing execution engines, optional rollout quantization, fully asynchronous rollouts, and rapidly evolving kernels. Our work is complementary: rather than assuming mismatch can be removed everywhere, we propose an algorithmic correction that remains robust when such mismatch persists, providing a unified importance-ratio formulation that tolerates heterogeneous sources of off-policy deviation. Although those methods are effective, they only focus on a specific problem and current RL training needs to combine the techniques and tune the parameters separately. The algorithm becomes even more complicated in fully-asynchronous settings (Fu et al., 2025) since there are three importance ratios. Hence, our work develops a unified importance ratio instead of dividing the importance ratio into several parts to solve general off-policy problems.
@@ 2 Related Works
Perturbation-based training is often motivated as a form of local averaging (randomized smoothing): optimizing the expected objective under small input/latent disturbances can reduce sensitivity to sharp, brittle regions of the loss landscape and improve robustness under distribution shift. This perspective has been used across a range of learning settings to stabilize optimization and enhance generalization.
A line of work has focused on enhancing model smoothness via perturbation-based sampling across diverse settings. For instance, Moreno-Barea et al. (2018); Yu et al. (2023) inject Gaussian noise to augment the training data and improve generalization. Certified robustness methods (Cohen et al., 2019; Salman et al., 2019; Lecuyer et al., 2019; Yang et al., 2020) pursue stability against adversarial attacks by introducing perturbations drawn from specific distributions during training. Beyond robustness, Li (2022); Pereira et al. (2021) demonstrate that perturbation injection can also improve performance across a range of tasks. Another related line of research focuses on diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal & Nichol, 2021; Saharia et al., 2022; Rombach et al., 2022), which model sample distributions as convolutions of Gaussian distributions and introduce Gaussian noise to the data during training. More recently, Shen & Meinshausen (2023); Hao et al. (2025) propose learnable perturbation mechanisms that adaptively enhance model performance. Related ideas also appear in diffusion models, where additional perturbations can mitigate train–test mismatch in iterative generation: Ning et al. (2023) reduce exposure bias by perturbing training inputs to better match the distribution encountered during sampling, while Li et al. (2023) alleviate exposure bias via sampling with shifted time steps.
@@ 3.1 Prior Approaches
where $|a_{i}|$ denotes the length of $a_{i}$, $A_{i}$ is the group advantage, and $\rho_{t}(\theta)$ is token-level importance ratio at $t$-th token
$\rho_{i,t}(\theta)=\frac{\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
@@
To balance the bias and variance, Group Sequence Policy Optimization (GSPO) (Zheng et al., 2025) proposes to use geometrical mean:
$\left(\frac{\pi^{\mathrm{train}}_{\theta}(a_{i}|x)}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i}|x)}\right)^{\frac{1}{|a_{i}|}}.$
@@
Since $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ and $\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}$ have a mismatch, to make algorithms robust to training-inference mismatch, a standard approach is to multiply by a correcting ratio (Yao et al., 2025; Liu et al., 2025):
@@
where $C$ is a threshold parameter, the gradient is masked when $\rho^{\prime}_{i,t}>C$, and the additional importance ratio
$\rho^{\prime}_{i,t}=\frac{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
This ratio can also be calculated in sequence level $\rho^{\prime}_{i}=\prod_{t}\rho^{\prime}_{i,t}$.
Bypass fail to stabilize training.
A direct idea to unify the importance ratio is directly replacing $\rho_{i}(\theta)$ with
$\frac{\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
This approach is a baseline in training-inference mismatch works (Yao et al., 2025; Jiacai et al., 2025), but it does not work effectively. Specifically, since the family class of $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ has system bias from the training policy $\pi_{\theta_{\mathrm{old}}}$, the policy gradients have noises and biases, and thus making the policy more unstable and brittle. Empirically, Bypass has blow up ratio tails during full training steps in Figure 1 (Right). Even with the same checkpoint and the same rollouts, the (log-)ratio envelope can flare up in the low-probability tail without additional mechanisms to control mismatch (Fig. 2 middle).
@@ 3.2 Our Approach
For layer $h\in[H]$ in the model, suppose the dimension of its input embedding is $d_{h}$, we add a Gaussian  [^2] We focus on Gaussian perturbations in experiments and analysis; extending to other zero-mean perturbations is an interesting direction. perturbation variable $\delta^{h}\sim\mathcal{N}(0,\sigma_{h}I_{d_{h}})$ to the updated training policy. Denoting the perturbation variable $\delta=[\delta^{1},\ldots,\delta^{H}]^{T}\sim\mathcal{N}(0,\sigma I_{d})$, where $\sigma$ is the std vector and $d=\sum_{h=1}^{H}d_{h}$, we define the perturbed policy
$\pi_{\theta,\sigma}(a|x,\delta)=\Pi_{t=1}^{|a|}\pi_{\theta,\sigma}(a_{t}|x,a_{<t},\delta_{t}),$
where $\delta_{t}$ matches the hidden-state tensor shape and is sampled independently across token positions. Then, the loss function is
$\mathcal{J}_{\mathrm{ALP}}(\theta,\sigma)=$ $\mathbb{E}_{x\sim d_{0},\{a_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x),\delta_{i}\sim\mathcal{N}(0,\sigma I_{d})}\bigg[\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{|a_{i}|}\min\Big\{\rho^{\mathrm{ALP}}_{i,t}(\theta)A_{i},\text{clip}\left(\rho^{\mathrm{ALP}}_{i,t}(\theta),1-\epsilon_{l},1+\epsilon_{h}\right)A_{i}\Big\}\bigg],$ (2)
@@ 3.3 Analysis: Robustness of ALP
Figure 2: Benefits of perturbation on smoothness: (Left) Toy simulation to show how perturbation smooths a sharp, spiky objective into a flatter surrogate, reducing sensitivity to local sharp maxima and promoting progress toward broader optima. (Middle, Right) Multi-turn, one-iteration controlled comparison. Starting from the same checkpoint and same batch of rollout samples, we perform $16$ off-policy updating steps without perturbation (middle) and with perturbation (right). Perturbation markedly shrinks the mismatch envelope, especially for low-probability tokens, thereby reducing extreme training–inference deviations.
The benefits of ALP can be stated in two aspects: both training-inference mismatch and smoothness. For simplicity, we conduct the theoretical analysis on one-layer perturbed policy. For each layer the perturbation $\delta$ is added to the embedding feature $x$, then ALP aims to learn a policy model with the expression as follows:
@@
For analytical convenience, we take the expectation over $\delta$. In practice, however, the algorithm samples $\delta$ only once per updating. This introduces a small discrepancy between the theoretical expectation and the single-sample approximation. We analyze the expectation form for clarity; in practice we use a single-sample approximation, which we find empirically sufficient in our setting.
Shrink the training-inference mismatch.
In the one-layer model, the training-inference mismatch can be formulated as
$\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\approx\pi_{\theta_{\mathrm{old}}}(a|x+\zeta),$
where the mismatch $\zeta$ is a zero-mean random variable.
For the original non-perturb method, taking the second-order Taylor expansion on KL distance, we have
$\quad\mathbb{E}_{x\sim d_{0}}\mathrm{KL}\left(\pi_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\approx\frac{1}{2}\zeta^{T}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}\zeta,$
which implies that
$\mathbb{E}_{\zeta}\mathbb{E}_{x\sim d_{0}}\mathrm{KL}\left(\pi_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\approx\frac{1}{2}\mathrm{Tr}\left(\mathbb{E}\zeta\zeta^{T}\cdot\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}\right).$ (4)
This approximation holds since the mismatch bias $\zeta$ is sufficiently small. Even with the small $\mathbb{E}\|\zeta\|_{2}^{2}$, as the eigenvalues of $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}$ can not be restricted during training process, a small training-inference mismatch is not guaranteed.
Now in ALP, with a proper perturbation variance $\sigma^{2}$, we could control the KL distance above effectively
@@
This theorem shows a trade-off on the scale of perturbation. Only when the perturbation std is small enough to stay close to the original distribution and large enough to cover the system noise norm $\|\zeta\|$, the KL-divergence between train and inference policy can be controlled effectively. The formal theorem is deferred to Theorem 3.
Moreover, TRPO shows the training is stabilized only when the distance between behavior and updated policy $\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\pi^{\mathrm{train}}_{\theta})$ is upper bounded by $\epsilon$ and the policy improvement is negatively affected by $\epsilon$ (Schulman et al., 2015). The distance is upper bounded by
$\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\widetilde{\pi}_{\theta})\leq\underbrace{\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\widetilde{\pi}_{\theta_{\mathrm{old}}})}_{\mathrm{Mismatch~Term}}+\underbrace{\mathrm{D}_{TV}(\widetilde{\pi}_{\theta_{\mathrm{old}}},\widetilde{\pi}_{\theta})}_{\mathrm{Controllable~Term}},$
where the first term is related to training-inference mismatch and the second term pertains to the training engine that is controllable and degrades to 0 when each batch has only one update. Hence, the instability mainly comes from Mismatch Term.
Without perturbation, as the mismatch between $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ can not be controlled from Equation (4), the condition $\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\pi_{\theta})\leq\epsilon^{\prime}$ in TRPO can not be easily satisfied. However, in ALP, Mismatch Term can be controlled by Theorem 3. Therefore, the updated policy $\widetilde{\pi}_{\theta}$ is more likely to fall into the trust region and guarantee training stability. In Figure 2 (Middle/Right), the quantile envelope shrinkage further validates this theory.
@@
On another side, adding perturbation can benefit the robustness of policy distribution, which implies that the smoothness can be smaller than the original algorithm. Without loss of generality, we neglect clipping and consider the standard reinforce loss and the perturbation loss:
$J(\theta)$ $=\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\left[\frac{\pi^{\mathrm{train}}_{\theta}(a|x)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot A(a)\right],$
$J_{\mathrm{ALP}}(\theta,\sigma)$ $=\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x),\delta\sim N(0,I)}\left[\frac{\pi^{\mathrm{train}}_{\theta}(a|x+\delta)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot A(a)\right].$
The main result is as follows (See the detailed proofs in Appendix C):
Theorem 2.
Define $\mathcal{I}(x,\theta)=\|\nabla^{2}_{\theta}J(\theta)\|_{2},\quad\widetilde{\mathcal{I}}(x,\theta)=\|\nabla^{2}_{\theta}J_{\mathrm{ALP}}(\theta,\sigma)\|_{2}$. If there exits a smallest constant $0<c<1$ such that
This result demonstrate that when the loss function is unsmooth, i.e., the curvature measured by the Hessian spectral norm $\mathcal{I}(x,\theta)$ varies substantially across different input points $x$, ALP can achieve better smoothness than the original non-perturbed method.
Empirical analysis.
The theories are demonstrated by empirical diagnostics. First, as shown in Figure 2 (Left), by introducing Gaussian perturbations, these high-curvature regions are effectively smoothed out. Thus, perturbation transforms a sharp objective into a locally smoothed surrogate, preventing optimization from collapsing onto spiky maxima and favoring flatter regions that are less sensitive to small distribution shifts. Additionally, in a controlled one-step intervention from the same checkpoint (Figure 2, middle/right), adding perturbation substantially narrows the conditional quantile envelope of the log-ratio, especially for low-probability tokens that dominate tail risk. This indicates that ALP also suppresses extreme training-inference deviations in a single update step, increasing the likelihood that updates remain within a trust region around the rollout distribution (Schulman et al., 2015).

### 29 Apr 2026 v2

@@ Abstract
Off-policy problems such as policy staleness and training–inference mismatch have become a major bottleneck for training stability and further exploration in LLM RL. The distribution gap between the inference and updated policies grows because of the techniques to enhance inference efficiency, leading to heavy-tailed importance ratios. Heavy-tailed ratios arise when the policy is locally sharp, which further inflates gradients and can push updates outside the trust region. To address this, we propose Adaptive Layerwise Perturbation (ALP), which injects small learnable perturbations into the input hidden states of each layer during updates and uses the resulting perturbed policy as the numerator of the importance ratio against the unchanged inference policy in the objective. Intuitively, by adding controlled noise to intermediate representations, ALP prevents the updated policy from deviating too sharply from the inference policy and enlarges the policy family to cover inference-time mismatch noise. Hence, the flattened distribution can naturally tighten the gap between the updated and inference policies and reduce the tail of importance ratios, thus maintaining training stability. This is further validated empirically. Experiments on single-turn math and multi-turn tool-integrated reasoning tasks show that ALP not only improves final performance, but also avoids blow-up in the importance-ratio tail and KL spikes during iterative training, along with boosted exploration. Ablations show that representation-level perturbations across all layers are most effective, substantially outperforming partial-layer and logits-only variants. Codes and training recipes are available at https://github.com/Chenluye99/Adaptive-Layerwise-Perturbation.
Figure 1: Left: ALP injects small layerwise perturbations during training to cover training-inference bias. Right: compared with Bypass over $1840$ steps, ALP yields a smoother policy and a tighter importance-ratio envelope, where lines P22, P9898 represent the 22th and $98$th quantile of the ratio.
@@ 1 Introduction
A prominent line of work stabilizes training by modifying importance ratios. One approach changes the aggregation way of importance ratios in the objective. It replaces the ratio at each token index in GRPO (Shao et al., 2024) with the multiplication of ratios in the whole trajectory (Zheng et al., 2025), but this does not consider training-inference mismatch. Another approach targets on training-inference mismatch. For example, Bypass is a baseline mentioned in (Jiacai et al., 2025; Yao et al., 2025) that uses the ratio of the updated and rollout policies. TIS/MIS (Yao et al., 2025; Jiacai et al., 2025) further correct mismatch by multiplying an auxiliary ratio of the proxy policy (an old policy acting as the anchor) and the rollout policy. Then, they truncate or mask tokens whose auxiliary ratio exceeds a threshold. Although effective at preventing catastrophic collapse, these methods introduce practical limitations: they split off-policy effects into two separately truncated ratios, which can over-truncate updates that are otherwise within a valid trust region, thereby inducing additional bias and leading to early plateau. We defer a detailed discussion of related work to Section 1.1.
@@
We argue that off-policy instability in LLM RL is not solely a bookkeeping issue of “which importance ratio to use”, but also a geometry issue: the noisy update may push the policy toward sharp, brittle regions where small distribution shifts (from staleness or inference mismatch) cause large changes in action probabilities. This motivates a perturbation-sampling perspective from model smoothness and distributional learning approaches (Shen and Meinshausen, 2023; Hao et al., 2025): when the dominant failure mode is system-induced noise and staleness, a natural defense is fighting noise with structured noise.
@@
Specifically, we propose Adaptive Layerwise Perturbation (ALP) by injecting a learnable Gaussian perturbation $\delta\sim\mathcal{N}(0,\sigma^{2}I)$ [^1] The perturbation distribution is not restricted to Gaussian; it only needs to have zero mean. into the input hidden states across layers during policy updates and producing a perturbed training distribution $\pi^{\mathrm{train}}_{\theta,\sigma}$ (Figure 1, left). We optimize the objective using a single unified ratio of $\pi^{\mathrm{train}}_{\theta,\sigma}$ and the unperturbed inference policy; i.e., perturbation is applied only to the numerator during training.
@@
Intuitively, ALP can smooth the local optimization landscape to suppress heavy-tailed ratio excursions and enlarge the training distribution family. This is also evidenced by the full-training comparison in Figure 1 (right). Starting from the same base model, the training policy without perturbation (Bypass) gradually becomes sharp and brittle, and the importance-ratio tail eventually explodes. In contrast, the training policy with perturbation (ALP) stays stable while controlling the mismatch between the training and inference systems, thus tightening the ratio-quantile envelope. Beyond stability, by enlarging effective support and preventing premature concentration on brittle modes, ALP also encourages exploration, particularly in multi-turn settings where compounding errors can reduce coverage.
@@
    Experiments: stability, performance, and ablations. Across single-turn and multi-turn agentic reasoning tasks, ALP consistently outperforms MIS and Bypass by keeping more stable KL divergence and avoiding early plateau. We further show that ALP reduces off-policy mismatch and exhibits more stable optimization dynamics, consistent with improved exploration and robustness. Ablations further reveal that perturbing all layers is most effective, and in partial settings, perturbations closer to lower layers tend to perform better than those restricted to upper layers.
1.1 Related Works
@@
Perturbation-based training is often motivated as a form of local averaging: optimizing the expected objective under small input/latent disturbances can reduce sensitivity to sharp, brittle regions of the loss landscape and improve robustness under distribution shift. This perspective has been used across a range of learning settings to stabilize optimization and enhance generalization. A line of work has focused on enhancing model smoothness via perturbation-based sampling across diverse settings. For instance, Moreno-Barea et al. (2018); Yu et al. (2023) inject Gaussian noise to augment the training data and improve generalization. Certified robustness methods (Cohen et al., 2019; Salman et al., 2019; Lecuyer et al., 2019; Yang et al., 2020) pursue stability against adversarial attacks by introducing perturbations drawn from specific distributions during training. Beyond robustness, Li (2022); Pereira et al. (2021) demonstrate that perturbation injection can also improve performance across a range of tasks. Another related line of research focuses on diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal and Nichol, 2021; Saharia et al., 2022; Rombach et al., 2022), which model sample distributions as convolutions of Gaussian distributions and introduce Gaussian noise to the data during training. More recently, Shen and Meinshausen (2023); Hao et al. (2025) propose learnable perturbation mechanisms that adaptively enhance model performance. Related ideas also appear in diffusion models, where additional perturbations can mitigate train–test mismatch in iterative generation: Ning et al. (2023) reduce exposure bias by perturbing training inputs to better match the distribution encountered during sampling, while Li et al. (2023) alleviate exposure bias via sampling with shifted time steps.
@@ 2.1 Prior Approaches
where $|a_{i}|$ denotes the length of $a_{i}$, $A_{i}$ is the group advantage, and $\rho_{i,t}(\theta)$ is the token-level importance ratio at the $t$-th token: $\rho_{i,t}(\theta)=\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})/\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t}).$
@@
To balance the bias and variance, Group Sequence Policy Optimization (GSPO) (Zheng et al., 2025) proposes using the geometric mean: $(\pi^{\mathrm{train}}_{\theta}(a_{i}|x)/\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i}|x))^{1/|a_{i}|}.$
@@
Since $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ and $\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}$ have a mismatch, to make algorithms robust to this mismatch, a standard approach is to multiply by a correcting ratio (Yao et al., 2025; Liu et al., 2025):
@@
where $C$ is a threshold parameter, the gradient is masked when $\rho^{\prime}_{i,t}>C$, and the additional importance ratio $\rho^{\prime}_{i,t}=\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})/\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t}).$ This ratio can also be calculated in sequence level $\rho^{\prime}_{i}=\prod_{t}\rho^{\prime}_{i,t}$.
Bypass fails to stabilize training.
A direct way to unify the importance ratio is to replace $\rho_{i}(\theta)$ with $\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})/\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t}).$ This approach is a baseline in training–inference mismatch work (Yao et al., 2025; Jiacai et al., 2025), but it does not work effectively. Specifically, since the family class of $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ has system bias relative to the training policy $\pi_{\theta_{\mathrm{old}}}$, the policy gradients contain noise and bias, making the policy more unstable and brittle. Empirically, Bypass exhibits blow-up in the ratio tail during full training in Figure 1 (right). Even with the same checkpoint and the same rollouts, the (log-)ratio envelope can flare up in the low-probability tail without additional mechanisms to control mismatch (Fig. 2, middle).
@@ 2.2 Our Approach
For layer $h\in[H]$ in the model, suppose the dimension of its input embedding is $d_{h}$. We add a Gaussianperturbation variable $\delta^{h}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d_{h}})$ to the updated training policy. Denoting the perturbation variable $\delta=[\delta^{1},\ldots,\delta^{H}]^{T}\sim\mathcal{N}(0,\mathrm{diag}(\sigma^{2}))$, where $\sigma$ is the std vector and $d=\sum_{h=1}^{H}d_{h}$, we define the perturbed policy $\pi_{\theta,\sigma}(a|x,\delta)=\prod_{t=1}^{|a|}\pi_{\theta,\sigma}(a_{t}|x,a_{<t},\delta_{t}),$ where $\delta_{t}$ matches the hidden-state tensor shape and is sampled independently across token positions. Then, the loss function is
$\mathcal{J}_{\mathrm{ALP}}(\theta,\sigma)=$ $\mathbb{E}_{x\sim d_{0},\{a_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x),\delta_{i}\sim\mathcal{N}(0,\mathrm{diag}(\sigma^{2}))}\bigg[\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{|a_{i}|}\min\Big\{\rho^{\mathrm{ALP}}_{i,t}(\theta)A_{i},\text{clip}\left(\rho^{\mathrm{ALP}}_{i,t}(\theta),1-\epsilon_{l},1+\epsilon_{h}\right)A_{i}\Big\}\bigg],$ (2)
@@ 2.3 Analysis: Robustness of ALP
Figure 2: Effects of perturbation. Left: perturbation smooths a sharp objective into a flatter surrogate. Middle/right: in a controlled multi-turn comparison from the same checkpoint and rollout batch, perturbation shrinks the mismatch envelope, especially in the low-probability tail.
ALP helps in two ways: it reduces training-inference mismatch and smooths the local optimization geometry. For clarity, we analyze a one-layer perturbed policy
@@
where the expectation is used only for analysis; the algorithm uses a single-sample approximation in practice. Formal statements and proofs are deferred to Appendix C and Appendix D.
Controlling training-inference mismatch.
We model the inference-time policy as the training-time policy with small noises on the input of each layer: $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\approx\pi_{\theta_{\mathrm{old}}}(a|x+\zeta),$ where $\zeta$ is a zero-mean system-induced bias. Without perturbation, even a small $\zeta$ can be amplified by a sharp local policy geometry, leading to unstable importance ratios. ALP mitigates this effect by enlarging the effective policy family around the training distribution and covering mismatch noise with perturbation.
@@
The formal one is deferred to Theorem 3. The theorem shows the key trade-off: $\sigma$ should be large enough to cover mismatch noise but small enough to preserve the original policy. In turn, the updated policy is more likely to stay close to the rollout distribution and remain within a trust region (Schulman et al., 2015).
@@
Beyond mismatch correction, perturbation acts as local averaging: instead of optimizing a sharp objective at a single representation, ALP optimizes an averaged neighborhood around it. This suppresses brittle local curvature and reduces sensitivity to spiky optima.
Theorem 2 (Informal).
Under mild regularity conditions, perturbation reduces the effective curvature of the surrogate objective, yielding a smoother optimization landscape than the non-perturbed objective.
We provide the formal theorem in Theorem 4. This result formalizes the intuition that ALP discourages optimization from collapsing onto sharp, fragile regions. Figure 2 provides empirical support: the toy example illustrates smoothing, while the controlled comparison shows a visibly tighter log-ratio envelope after applying perturbation. Additionally, in a controlled one-step intervention from the same checkpoint (Figure 2, middle/right), adding perturbation substantially narrows the conditional quantile envelope of the log-ratio, especially for low-probability tokens that dominate tail risk. This indicates that ALP also suppresses extreme training-inference deviations in a single update step, increasing the likelihood that updates remain within a trust region around the rollout distribution (Schulman et al., 2015).